The perimeter of a rectangle is 28 meters.If the length of the rectangle is (2 + x 𝑥) and the width is 3, then what does x equal?

Answer by fractalier(6550)

The perimeter of a rectangle is 28 meters. If the length of the rectangle is (2 + x) and the width is 3, then what does x equal?
Perimeter P = 2L + 2W = 28...we have
2(2+x) + 2(3) = 28
4 + 2x + 6 = 28
2x + 10 = 28
2x = 18
x = 9 meters